OpenStudy (anonymous):

Write an equation of a line in slope intercept form that is parallel to the line x + 4y = 6 and passes through (-8, 5).

OpenStudy (anonymous):

x+4y=6 4y=-x+6 y=(-1/4)x+(6/4) here,slope of line is -1/4. so,by problem, (y-5)=-1/4(x+8) 4y-20=-x-8 4y+x=-8+20 4y+x=12 this is the required equation.

OpenStudy (anonymous):

thank you veryy veryyy much!

OpenStudy (anonymous):

u r most welcme!!!